OpenStudy (anonymous):

Solve by elimination 2x – 5y = 1 3x – 4y = -2

OpenStudy (lanhikari22):

(x3) 6x - 15y = 3 (x -2) -6x + 8y = 4 add the equations together -7y = 7 y = -7/7 = -1

OpenStudy (lanhikari22):

also, 2x -5(-1) = 1 so 2x + 5 = 1 2x = -4 x = -2

OpenStudy (texaschic101):

2x - 5y = 1 --->(3) 3x - 4y = -2 -->(-2) ------------------ 6x - 15y = 3 (result of multiplying by 3) -6x + 8y = 4 (result of multiplying by -2) --------------------add -7y = 7 y = -1 2x - 5y = 1 2x - 5(-1) = 1 2x + 5 = 1 2x = 1 - 5 2x = -4 x = -2 solution is (-2,-1) yep...I agree with LanHikari :)

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

no problem :)